A body of weight 6N rests on a inclined plane of 30° to the horizontal.What is the force keeping it from sliding down the plane and the coefficient of friction


1
Expert's answer
2021-05-12T08:37:36-0400
"F=6\sin{30}=3\ N\\\mu=\frac{3}{6\cos{30}}=0.577"